### v1

@@ 6.2 Simulation Environment
The router simulator models an online setting with $G$ GPU workers, each supporting up to $B$ concurrent requests (batch size). Key components include:
@@
where $C$ is a fixed overhead constant and $t_{\ell}$ is the per-token generation latency. During each step, all active requests generate one token. A request $i$ completes when $\tau_{i}\geq o_{i}$.
Online Revelation.
The simulator maintains a revelation target $R$; whenever $|\text{wait\_q}|<R$, additional requests are revealed from the undiscovered queue. This models realistic online arrival patterns. The default configuration of our experiment is $G=32$ GPUs, $B=72$ batch size, $R=128$ revelation target.
Routing Strategies.
We evaluate four scheduling policies: (i) FCFS, (ii) Join-Shortest-Queue (JSQ), (iii) BF-IO ($H=0$), (iv) BF-IO ($H=20$)
@@ 6.3 Evaluation Metrics
$E=\int_{0}^{T}P\bigl(\mathrm{mfu}(t)\bigr)\,dt$ (10)
where $\mathrm{mfu}(t)$ denotes the Model FLOPs Utilization (mfu) at time tt—the ratio of observed computational throughput to the GPU’s theoretical peak FLOPs capacity. Therefore, the instant mfu at time $t$ is positively related to the instant throughput at $t$. [ozcan2025quantifying] validates instantaneous power is a sublinear function of utilization:
with idle power $P_{\mathrm{idle}}=100$W, peak power $P_{\mathrm{max}}=400$W, saturation threshold $\mathrm{mfu}_{\mathrm{sat}}=0.45$, and scaling exponent $\gamma=0.7$. The sublinear exponent $\gamma<1$ captures a critical inefficiency: GPUs draw substantial power even at low utilization. Under barrier synchronization, workers waiting at the barrier operate near $P_{\mathrm{idle}}$ while contributing zero throughput; reducing imbalance converts this idle power draw into useful token generation, directly improving energy efficiency. Details of MFU computation and parameter calibration are provided in Appendix D.
@@ 6.4 Results
Figure 6 visualizes the per-worker token load trajectories over approximately 4000 decode steps for four scheduling strategies on a 16-GPU configuration. Each colored line represents the total token load $L_{g}(k)$ on a single worker across time.
Figure 6: Per-worker token load trajectories under four scheduling strategies. Top-left: FCFS exhibits persistent load spread between 60k–110k tokens, with substantial inter-worker variance throughout execution. Top-right: JSQ shows even greater imbalance, with peak loads exceeding 120k while other workers remain around 80k. Bottom-left:BF-IO ($H=0$) achieves significantly tighter load clustering around 80k tokens. Bottom-right: BF-IO ($H=20$) produces the most balanced distribution, with all workers maintaining nearly identical loads throughout the simulation.
The results reveal stark differences in load balancing effectiveness:
  • FCFS maintains a persistent load gap of approximately 40k–50k tokens between the heaviest and lightest workers. This spread translates directly into barrier-induced idle time, as lightly loaded workers must wait for the straggler at each decode step.
  • JSQ, despite being workload-aware, exhibits even worse imbalance than FCFS. This counterintuitive result aligns with the theoretical analysis: JSQ uses queue length (number of requests) as a proxy for workload, which poorly approximates actual token load when request sizes vary significantly. Heavy requests accumulate on certain workers, creating persistent stragglers.
  • BF-IO ($H=0$) dramatically reduces the load spread to approximately 5k–10k tokens. Even without lookahead information, myopically minimizing instantaneous imbalance at each assignment step yields substantial improvements over size-agnostic policies.
  • BF-IO ($H=20$) achieves the tightest load balance, with worker loads nearly indistinguishable throughout execution. The 80-step lookahead enables the scheduler to anticipate near-term completions and avoid assignments that would create transient imbalance.
@@
Table 1 summarizes the performance metrics across all evaluated strategies. We report average imbalance, throughput, and TPOT, each averaged over multiple simulation runs with different random seeds.
Table 1: Performance comparison of scheduling strategies.
@@
[…]
FCFS 27.9 8.00 1.42 396
JSQ 28.2 7.99 1.42 396
BF-IO ($H=0$) 2.92 9.03 1.26 386
BF-IO ($H=20$) 1.65 9.13 1.25 383
Table 1 summarizes the performance metrics across all evaluated strategies. We report average imbalance, throughput, TPOT, and energy consumption, each averaged over multiple simulation runs with different random seeds.
[…]
Figure 7: Average GPU power consumption over time for FCFS and BF-IO ($H=20$).
Figure 7 illustrates the average GPU power consumption over time for the FCFS and BF-IO $(H=20)$ scheduling strategies. The BF-IO strategy maintains a consistently high power draw around 395–400W throughout the experiment, whereas the Default strategy exhibits significant fluctuations between 330W and 380W.
Although BF-IO operates at higher instantaneous power levels, it achieves lower total energy consumption (383 kJ vs. 396 kJ). This counterintuitive result can be attributed to improved load balancing: BF-IO keeps all GPUs uniformly utilized, avoiding idle periods and reducing the overall execution time. In contrast, the power dips observed in the Default strategy indicate periods of underutilization caused by load imbalance across GPUs. These results demonstrate that better workload distribution not only improves throughput and latency but also leads to improved energy efficiency.
@@ 7.1 Summary of contributions
This work develops a universal load-balancing principle for barrier-synchronized systems with sticky assignment decisions. We formalize the principle as BF-IO, which selects assignments by minimizing a short-horizon prediction of the imbalance metric induced by the slowest worker via an integer optimization. We establish worst-case, scale-sensitive guarantees for BF-IO under adversarial arrivals in the LLM decode model and extend the analysis to a broader class of non-decreasing workload drift processes. On the empirical side, extensive numerical experiments on both public benchmarks and proprietary industrial traces demonstrate that BF-IO substantially reduces imbalance, improves throughput, and lowers energy consumption relative to widely deployed baselines.

### v2

@@ 6.2 Simulation Environment
The router simulator models an online setting with $G=256$ A100 GPU workers, each supporting up to $B=72$ concurrent requests (batch size). Key components include:
@@
where $C=9.775\cdot 10^{-3}$ is a fixed overhead constant and $t_{\ell}=1.005\cdot 10^{-7}$ is the per-token generation latency. The values of $C$ and $t_{\ell}$ are derived from a minimum mean square error regression on real-world traces. During each step, all active requests generate one token. A request $i$ completes when $\tau_{i}\geq o_{i}$.
Routing Policies.
We evaluate four routing policies: (i) FCFS, (ii) Join-Shortest-Queue (JSQ), (iii) BF-IO ($H=0$), (iv) BF-IO ($H=20$), (vi) BF-IO ($H=40$), (vii) BF-IO ($H=60$), (viii) BF-IO ($H=80$), (ix) BF-IO ($H=100$).
@@ 6.3 Evaluation Metrics
As is mentioned in Section 5.2, the total energy consumption (in Joules) is the time integral of instantaneous GPU power (Equation (6)), where the instantaneous GPU power can be expressed by Equation (7). Details of MFU computation and parameter calibration are provided in Appendix D.
@@ 6.4 Results
We present quantitative comparisons across routing policies, supported by visualization of load dynamics and power consumption patterns.
@@
Figure 7 presents per-worker load trajectories over the processing interval. To maintain visual clarity with $G=256$ workers, we display a random sample of 16 GPUs. Under the baseline policies FCFS and JSQ, worker loads exhibit pronounced fluctuations ranging from 10M to 35M tokens during stable decoding, with high-frequency oscillations reflecting severe and persistent imbalance across the cluster. BF-IO ($H=0$) substantially compresses the load spread to approximately 12M–23M tokens. With a 40-step lookahead, BF-IO ($H=40$) achieves markedly tighter clustering (16M–17M tokens) despite the challenging workload characteristics, demonstrating the robustness of the lookahead mechanism under high-variance conditions.
Figure 7: Per-worker token load trajectories under four routing policies (16 randomly sampled workers from $G=256$). FCFS and JSQ exhibit extreme fluctuations (10M–35M tokens), whereas BF-IO policies maintain substantially tighter distributions, with BF-IO ($H=40$) achieving near-uniform loads around 16M–17M tokens.
Power Consumption Analysis.
Figure 8 illustrates the temporal dynamics of average GPU power consumption, revealing how routing policies affect both aggregate energy and instantaneous power stability.
Figure 8: Average GPU power consumption over time. BF-IO sustains consistently high power draw (395–400 W), indicating near-full utilization, and completes the workload faster. The baseline FCFS policy exhibits substantial fluctuations (270–360 W) due to load imbalance. Despite higher instantaneous power, BF-IO achieves 30% lower total energy consumption (20.9 MJ vs. 29.1 MJ) by eliminating idle time.
During stable decoding (excluding initial ramp-up), BF-IO maintains consistently high instantaneous power (395–400 W), indicating balanced utilization across all devices. For reference, the NVIDIA A100 GPU has a maximum power draw of $P_{\mathrm{max}}=400$ W (Appendix C.4), implying that BF-IO operates near full computational capacity throughout execution. In contrast, the baseline FCFS policy oscillates between 270–360 W, reflecting periods in which subsets of GPUs remain underutilized due to uneven workload distribution.
Crucially, although BF-IO operates at higher instantaneous power, it achieves substantially lower total energy consumption (20.9 MJ vs. 29.1 MJ for FCFS). This apparent paradox resolves by recognizing that total energy is the time integral of instantaneous power: by maintaining all GPUs at high utilization, BF-IO completes the workload faster, reducing total execution time and thereby the integral. In effect, balanced scheduling eliminates idle power—energy expended without productive computation—yielding significant net savings.
@@
Table 1 summarizes performance across routing policies. BF-IO consistently outperforms baselines across all metrics for every choice of lookahead window $H$ tested.
Table 1: Performance comparison on LongBench. Arrows indicate preferred direction. Best values in each column are highlighted.
@@
($\times 10^{8}$) ($10^{3}$ tok/s) (s/tok) (MJ)
FCFS 37.0 5.47 2.97 29.1
JSQ 28.5 6.15 2.64 27.8
BF-IO ($H=0$) 9.94 8.55 1.93 23.4
BF-IO ($H=20$) 3.03 10.1 1.70 21.3
BF-IO ($H=40$) 2.49 10.5 1.67 20.6
BF-IO ($H=60$) 2.44 10.4 1.66 20.9
BF-IO ($H=80$) 2.43 10.4 1.70 21.0
BF-IO ($H=100$) 2.51 10.4 1.67 20.9
The optimal lookahead window is $H=40$. Relative to FCFS, BF-IO ($H=40$) reduces average imbalance by a factor of $15\times$, improves throughput by 92%, reduces time-per-output-token (TPOT) by 44%, and lowers energy consumption by 29%. These gains are substantial and consistent across metrics.
Comparison across BF-IO variants reveals a tradeoff in selecting $H$ (Figure 9). When $H$ is too small, the policy optimizes myopically, balancing loads only over the immediate future and missing opportunities to anticipate near-term completions. When $H$ is too large, the optimization horizon extends beyond the range of reliable prediction—since arrivals are online and unpredictable, the objective becomes biased toward stale information. Additionally, larger $H$ incurs greater computational overhead per decision. Empirically, performance saturates around $H=40$; further increases yield no benefit and marginally degrade some metrics.
Figure 9: Effect of lookahead horizon $H$ on key performance metrics. All metrics improve rapidly as $H$ increases from 0 to 40, then plateau. The optimal operating point balances prediction reliability against computational overhead; beyond $H\approx 40$, additional lookahead provides no further benefit.
6.5 Ablation Study
Scalability with Cluster Size.
To understand how BF-IO scales with increasing computational resources, we conduct an ablation study varying the number of GPU workers $G$ from 16 to 224 while keeping the workload fixed (Longbench dataset). Figures 10 and 11 present the scaling behavior of BF-IO compared to FCFS.
Figure 10: Scalability analysis with varying cluster size $G$. Left: Average imbalance grows super-linearly under FCFS but remains bounded under BF-IO. Right: BF-IO achieves near-linear throughput scaling while FCFS exhibits diminishing returns.
The left panel of Figure 10 reveals that the average imbalance under FCFS grows super-linearly with cluster size, increasing from approximately $3\times 10^{8}$ at $G=16$ to over $3.5\times 10^{9}$ at G=224G=224—a $12\times$ increase for a $7\times$ scale-up in resources. In contrast, BF-IO maintains imbalance below $3\times 10^{8}$ across all configurations, demonstrating that workload-aware scheduling effectively contains the coordination overhead even at scale. At $G=224$, BF-IO achieves $14\times$ lower imbalance than FCFS.
The right panel shows that throughput scaling exhibits fundamentally different characteristics between the two strategies. BF-IO achieves near-linear scaling, increasing from approximately 1k tokens/s at $G=16$ to over 9k tokens/s at $G=224$. Meanwhile, FCFS throughput grows sub-linearly, reaching only about 4.5k tokens/s at the largest scale—less than half of BF-IO’s throughput. This divergence stems directly from the imbalance growth: as more workers are added under FCFS, the probability of severe load skew increases, causing stragglers that bottleneck overall progress.
Figure 11: Energy consumption scaling with cluster size. BF-IO (blue) achieves consistently lower energy than FCFS (orange), with energy reduction (green, right axis) increasing from 12% at $G=16$ to 30% at $G=224$.
Figure 11 highlights the operational efficiency gains of BF-IO. While FCFS energy consumption remains relatively flat around 31–34 MJ across different cluster sizes, BF-IO energy decreases steadily from approximately 29 MJ at $G=16$ to 22 MJ at $G=224$. The energy reduction percentage (green line) increases monotonically with cluster size, rising from 12% at small scales to 30% at $G=224$. This widening gap demonstrates that the benefits of intelligent scheduling compound at larger scales: not only does BF-IO complete workloads faster, but it also avoids the energy waste from idle GPUs waiting at synchronization barriers.
@@ 7.1 Summary of contributions
This work develops a universal load-balancing principle for barrier-synchronized systems with sticky assignment decisions. We formalize the principle as BF-IO, which selects assignments by minimizing a short-horizon prediction of the imbalance metric induced by the slowest worker via an integer optimization. We establish worst-case, scale-sensitive guarantees for BF-IO under adversarial arrivals in the LLM decode model and extend the analysis to a broader class of non-decreasing workload drift processes. Crucially, we prove that this imbalance reduction yields rigorous energy savings guarantees: as system scale grows, the guaranteed energy reduction approaches 52% for modern GPU hardware. This result establishes, to our knowledge, the first theoretical link between load balancing policy and provable energy efficiency in LLM serving. On the empirical side, extensive numerical experiments on both public benchmarks and proprietary industrial traces demonstrate that BF-IO substantially reduces imbalance, improves throughput, and lowers energy consumption relative to widely deployed baselines.
@@ C.3 Proof of Theorem 3
C.4 Proof of Theorem 4
Proof of Theorem 4.
Fix $\pi$ and $\mathcal{I}$. For each step $k$, recall $u_{g}(k)=L_{g}(k)/L_{g^{*}}(k)$. By (10),
$E(\pi;\mathcal{I})$ $=\kappa_{\mathrm{ATT}}\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}L_{g^{*}}(k)\sum_{g=1}^{G}\Bigl(P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}\Bigr)$
$=\kappa_{\mathrm{ATT}}\sum_{k,g}\Bigl(P_{\mathrm{idle}}L_{g^{*}}(k)+(P_{\mathrm{max}}-P_{\mathrm{idle}})L_{g^{*}}(k)u_{g}(k)\Bigr)\;$ (66)
$+\;\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})\sum_{k,g}L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr),$ (67)
where $\sum_{k,g}$ abbreviates $\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}$. Since $L_{g^{*}}(k)u_{g}(k)=L_{g}(k)$, the first sum in (67) becomes
$\kappa_{\mathrm{ATT}}\sum_{k,g}\Bigl(P_{\mathrm{idle}}L_{g^{*}}(k)+(P_{\mathrm{max}}-P_{\mathrm{idle}})L_{g}(k)\Bigr)$ $=\kappa_{\mathrm{ATT}}\sum_{k,g}\Bigl(P_{\mathrm{max}}L_{g}(k)+P_{\mathrm{idle}}(L_{g^{*}}(k)-L_{g}(k))\Bigr)$
$=\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I}),$ (68)
using (11) and (12). Substituting (68) into (67) yields the exact identity
$E(\pi;\mathcal{I})=\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I})+\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})\sum_{k,g}L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr).$ (69)
We next bound the last term in (69). Since $\gamma\in(0,1)$ and $u\in[0,1]$, we have $u^{\gamma}\geq u$, hence the last term is nonnegative. For an upper bound, concavity of $u^{\gamma}$ on $[0,1]$ implies the tangent inequality at $u=1$:
$u^{\gamma}\;\leq\;1+\gamma(u-1)\;=\;(1-\gamma)+\gamma u,$
so $u^{\gamma}-u\leq(1-\gamma)(1-u)$. Therefore,
$L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr)\;\leq\;(1-\gamma)\,L_{g^{*}}(k)\bigl(1-u_{g}(k)\bigr)=(1-\gamma)\bigl(L_{g^{*}}(k)-L_{g}(k)\bigr).$
Summing over $k,g$ and using (12),
[…]
Plugging (70) into (69) yields the sandwich bound
$\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I})\ \leq\ E(\pi;\mathcal{I})\ \leq\ \kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}C_{\gamma}\,\mathrm{ImbTot}(\pi;\mathcal{I}),$ (71)
where $C_{\gamma}$ is defined in (15).
Now apply (71) with $\pi=\pi_{0}$ and $\pi=\pi_{1}$. Using the upper bound for $E(\pi_{0};\mathcal{I})$ and the lower bound for $E(\pi_{1};\mathcal{I})$ gives
$E(\pi_{0};\mathcal{I})\ \leq\ \kappa_{\mathrm{ATT}}W(\mathcal{I})\Bigl(P_{\mathrm{max}}+C_{\gamma}\,\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})\Bigr),$ (72)
where we used $\mathrm{ImbTot}(\pi_{0};\mathcal{I})=\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})W(\mathcal{I})$. For the energy difference, use (69) and (70) to obtain
$E(\pi_{0};\mathcal{I})-E(\pi_{1};\mathcal{I})$ $=\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\bigl(\mathrm{ImbTot}(\pi_{0};\mathcal{I})-\mathrm{ImbTot}(\pi_{1};\mathcal{I})\bigr)$
$+\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})\bigl(X(\pi_{0})-X(\pi_{1})\bigr),$ (73)
where $X(\pi):=\sum_{k,g}L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr)$. By (70), $X(\pi_{0})\geq 0$ and $X(\pi_{1})\leq(1-\gamma)\mathrm{ImbTot}(\pi_{1};\mathcal{I})$. Hence
$E(\pi_{0};\mathcal{I})-E(\pi_{1};\mathcal{I})$ $\geq\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\bigl(\mathrm{ImbTot}(\pi_{0};\mathcal{I})-\mathrm{ImbTot}(\pi_{1};\mathcal{I})\bigr)$
$-\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})(1-\gamma)\,\mathrm{ImbTot}(\pi_{1};\mathcal{I})$
$=\kappa_{\mathrm{ATT}}\left(P_{\mathrm{idle}}\bigl(\mathrm{ImbTot}(\pi_{0};\mathcal{I})-\mathrm{ImbTot}(\pi_{1};\mathcal{I})\bigr)-D_{\gamma}\,\mathrm{ImbTot}(\pi_{1};\mathcal{I})\right),$ (74)
where $D_{\gamma}$ is defined in (15). Under the assumption (14), $\mathrm{ImbTot}(\pi_{1};\mathcal{I})\leq\mathrm{ImbTot}(\pi_{0};\mathcal{I})/\alpha$, so (74) implies
$E(\pi_{0};\mathcal{I})-E(\pi_{1};\mathcal{I})\ \geq\ \kappa_{\mathrm{ATT}}\mathrm{ImbTot}(\pi_{0};\mathcal{I})\left(P_{\mathrm{idle}}\Bigl(1-\frac{1}{\alpha}\Bigr)-\frac{D_{\gamma}}{\alpha}\right).$ (75)
Combining (75) with (72) and using $\mathbb{E}\bigl[\mathrm{ImbTot}(\pi_{0};\mathcal{I})\bigr]=\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})\mathbb{E}\bigl[W(\mathcal{I})\bigr]$ gives (16). Taking the expectation completes the proof. ∎
C.5 Proof of Inequality (17)
Proof of Inequality (17).
Under FCFS in the overloaded regime, the slot-level Markov recursion (36) (Section 5) is uniformly ergodic with unique stationary slot law $U\overset{d}{=}S_{\mathrm{cur}}+A$ where $S_{\mathrm{cur}}\stackrel{{\scriptstyle d}}{{=}}S$ and $A\sim\mathrm{Geo}(p)-1$ independent, hence
$\mu_{U}:=\mathbb{E}[U]=\mu_{s}+\frac{1-p}{p},\qquad\mathrm{Var}(U)=\sigma_{\mathrm{snap}}^{2}.$
At stationarity, the total system workload in a step equals $\sum_{g=1}^{G}\sum_{b=1}^{B}U_{g,b}$, so its expectation is $GB\mu_{U}$. Therefore the long-run average total workload per step under FCFS equals $GB\mu_{U}$ by uniform ergodicity and Cesàro averaging. On the other hand, Part 2 of Appendix C.2 establishes the lower bound
$\mathbb{E}\bigl[\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\bigr]\ \geq\ c^{\prime}\,G\,\sigma_{\mathrm{snap}}\sqrt{B\log G}$
for a universal constant $c^{\prime}>0$ and all $B$ sufficiently large. Dividing by $GB\mu_{U}$ yields
$\frac{\mathbb{E}\bigl[\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\bigr]}{GB\mu_{U}}\ \geq\ c^{\prime}\,\frac{\sigma_{\mathrm{snap}}}{\mu_{U}}\sqrt{\frac{\log G}{B}}.$
Interpreting $\eta_{\mathrm{sum}}(\textsf{FCFS};\mathcal{I})$ as the corresponding normalized imbalance level (cumulative or long-run averaged; these coincide under stationarity), we obtain (17) (absorbing constants into $\gtrsim$). ∎
@@ D.1 Energy and MFU Computation
D.2 Results on the Lighter-Load Dataset BurstGPT
We additionally evaluate our routing policy on BurstGPT [wang2024burstgpt], a lighter-load benchmark with shorter request lengths than LongBench. The prefill and decode length distributions are shown in Figure 12. Given the reduced per-request workload, 16 GPUs suffice to process this dataset.
Figure 12: Distribution of prefill (input) and decode (output) lengths in the BurstGPT dataset.
Load Trajectory Analysis.
Figure 13 presents per-worker load trajectories for the BurstGPT workload. Under FCFS, worker loads spread across 250k–400k tokens, with persistent inter-worker variance throughout execution. JSQ exhibits even greater imbalance, confirming that queue length is a poor proxy for actual token load when request sizes vary substantially. In contrast, BF-IO ($H=0$) achieves tighter clustering around 270k–320k tokens. With lookahead enabled, BF-IO ($H=20$) produces the most balanced distribution: all workers maintain nearly identical loads throughout execution, reducing average imbalance by a factor of $17\times$ compared to FCFS.
Figure 13: Per-worker token load trajectories under four routing policies (BurstGPT, $G=16$ GPUs). FCFS exhibits persistent load spread (250k–400k tokens); JSQ shows even greater imbalance due to its size-agnostic assignment. BF-IO ($H=0$) achieves substantially tighter clustering, and BF-IO ($H=20$) produces near-uniform loads across all workers throughout execution.
Power Consumption Analysis.
Figure 14 illustrates average GPU power consumption over time. The baseline FCFS policy oscillates between 330–380 W, reflecting periods in which subsets of GPUs remain underutilized. BF-IO ($H=20$) maintains more stable power draw at higher sustained utilization, completing the workload faster despite drawing similar peak power—consistent with the energy-saving mechanism established in the main text.
Figure 14: Average GPU power consumption over time (BurstGPT, $G=16$ GPUs). FCFS exhibits fluctuations (330–380 W) due to load imbalance, while BF-IO maintains higher sustained utilization and completes the workload in less time, yielding lower total energy consumption.
Quantitative Metrics Comparison.
Table 2 summarizes performance across routing policies. Even in this lighter-load regime, BF-IO ($H=20$) reduces average imbalance by $17\times$ relative to FCFS, improves throughput by 14%, reduces TPOT by 12%, and lowers energy consumption by 3.3%. The more modest energy savings compared to LongBench (3.3% vs. 29%) reflect the lower baseline imbalance in this easier workload; nevertheless, the improvements remain consistent and statistically significant across all metrics.
Notably, FCFS and JSQ exhibit nearly identical performance on this dataset, confirming that both policies are effectively size-agnostic: JSQ balances queue lengths rather than token loads, providing no advantage when request sizes vary.
Table 2: Performance comparison on BurstGPT ($G=16$ GPUs, $B=72$ batch size). Arrows indicate preferred direction.
[…]
($\times 10^{4}$) ($10^{4}$ tok/s) ($10^{-2}$ s/tok) (kJ)
FCFS 27.9 8.00 1.42 396
JSQ 28.2 7.99 1.42 396
BF-IO ($H=0$) 2.92 9.03 1.26 386
BF-IO ($H=20$) 1.65 9.13 1.25 383